OpenStudy (anonymous):

Which values are solutions to the inequality below √x ≥10 A)120 B)101 C)19 D)567 E)5 F)100

OpenStudy (ash2326):

\[\sqrt x \ge 10\] Since we have a sqrt, we could square both sides squaring both sides \[x \ge 100\] Now can you tell me the correct options

OpenStudy (anonymous):

F,D,B AND A?

OpenStudy (ash2326):

Yeah, you're correct:D

OpenStudy (anonymous):

THANX :DDDDDDD

OpenStudy (anonymous):